OpenStudy (lovedance11):

Subtract. 7 - 5 - - 16 16

Mehek (mehek14):

I just helped you with that ._.

OpenStudy (lovedance11):

I need to ask you is it 2/16

Mehek (mehek14):

yes

Mehek (mehek14):

simplify to 1/8

OpenStudy (lovedance11):

im doing an assignment and it wasnt one of the answer choices

Mehek (mehek14):

what are your choices

OpenStudy (welshfella):

7 - 5 = 2 so its 2/16 which simplifies to 1/8

OpenStudy (lovedance11):

1/8 1/4 12/32 3/4

Mehek (mehek14):

i told you 2/16 = 1/8

OpenStudy (lovedance11):

Thx everyone

OpenStudy (ninjaslice):

its 1/8 =) please medal me